# Bidirectional Prototype-Reward co-Evolution for Test-Time Adaptation of Vision-Language Models

### v1

@@ Abstract
Test-time adaptation (TTA) is crucial in maintaining Vision Language Models(VLMs)’ performance when facing real-world distribution shifts, particularly when the source data or target labels are inaccessible. Existing TTA methods rely on CLIP’s output probability distribution for feature evaluation, which can introduce biases under domain shifts. This misalignment may cause features to be misclassified due to text priors or incorrect textual associations. To address these limitations, we propose Bidirectional Prototype-Reward co-Evolution (BPRE), a novel TTA framework for VLMs that integrates feature quality assessment with prototype evolution through a synergistic feedback loop. BPRE first employs a Multi-Dimensional Quality-Aware Reward Module to evaluate feature quality and guide prototype refinement precisely. The continuous refinement of prototype quality through Prototype-Reward Interactive Evolution will subsequently enhance the computation of more robust Multi-Dimensional Quality-Aware Reward Scores. Through the bidirectional interaction, the precision of rewards and the evolution of prototypes mutually reinforce each other, forming a self-evolving cycle. Extensive experiments are conducted across 15 diverse recognition datasets encompassing natural distribution shifts and cross-dataset generalization scenarios. Results demonstrate that BPRE consistently achieves superior average performance compared to state-of-the-art methods across different model architectures, such as ResNet-50 and ViT-B/16. By emphasizing comprehensive feature evaluation and bidirectional knowledge refinement, BPRE advances VLM generalization capabilities, offering a new perspective on TTA.
1 Introduction
@@
Recent advances in Vision-Language Models (VLMs) [2, 38, 18] have opened new opportunities to integrate human language into various computer vision tasks. Despite the unprecedented zero-shot generalization capability achieved by VLMs such as CLIP [23], existing VLMs suffer from a large performance drop when encountering domain shifts, which is due to the inconsistent data distribution between the source domain and the target domain. However, in real-world applications, the inevitable domain discrepancies caused by dynamic and ever-changing environments make domain adaptation of VLMs an urgent problem.
To address this issue, some researchers have developed efficient adaptation methods [40, 45, 3] that fine-tune the VLMs using limited target domain samples. However, the requirement for target domain annotations remains a critical bottleneck, especially in dynamic environments with frequent distribution shifts [24, 36]. Recently, several studies have explored the test-time adaptation (TTA) for VLMs, aiming to adaptively optimize the model at a low cost during the test phase. Shu et al. [25] introduced test-time prompt tuning (TPT) that optimizes prompts for individual test samples to enhance out-of-distribution generalization. This framework was enhanced by DiffTPT [8] through diffusion-based augmentations, while Karmanov et al. [14] proposed a training-free approach using dynamic visual caches. More recently, Zhang et al. [37] developed a dual prototype evolving strategy that accumulates task-specific knowledge from both visual and textual modalities during testing.
Overall, most of the existing TTA works [40, 14, 37] have highlighted the importance of utilizing historical feature information to accumulate task-specific knowledge for each class by Caches and Prototypes. This approach aims to effectively construct accurate and discriminative representations of target classes, which are crucial for enhancing the model’s robustness to distribution shifts, as well as improving its efficiency and performance during the test-time phase. Current methods [14, 37] primarily rely on CLIP’s output probability distribution, derived from the fusion of Text Features and Image Features, to evaluate feature quality, inherently using the text modality as a reference. However, as shown in Fig. 2, distribution shifts during testing may cause the text encoder to misalign with the semantic space of the image modality, leading to biases in feature selection and evaluation. On one hand, text modality priors cause the model to favor features that strongly align with pre-trained textual descriptions, even when they lack sufficient visual information, while visually rich samples with weaker textual representations may be misclassified (Fig. 2(a)). On the other hand, misalignment of textual descriptions can introduce feature evaluation biases, where CLIP may incorrectly associate generalized textual descriptions (e.g.,“picture”) with certain image features, leading the model to mistakenly consider them high-quality samples while overlooking their true semantics, thereby affecting feature selection accuracy (Fig. 2(b)).
Based on these insights, we propose a novel test-time VLM adaptation framework: Bidirectional Prototype-Reward co-Evolution (BPRE), which iteratively refines task-specific knowledge, as illustrated in Fig. 3. Unlike existing approaches, BPRE introduces the Multi-Dimensional Quality-Aware Reward Module, an evaluation method that integrates intrinsic Image Feature properties. By leveraging visual information while incorporating Semantic Similarity and Feature Diversity as key criteria, this module provides a more comprehensive assessment of Image Feature quality, reducing excessive dependence on the text modality. Semantic Similarity ensures selected features align with the task’s semantic space, mitigating errors from textual biases, while Feature Diversity enhances coverage across varied data distributions, preventing overfitting and improving prototype selection during test-time adaptation. Additionally, the Prototype-Reward Interactive Evolution module adaptively updates historical features and Prototypes using Reward scores. This bidirectional reinforcement loop allows refined prototypes to enhance reward estimation, forming a self-reinforcing optimization process that boosts robustness against noise and facilitates interpretable cross-modal learning, yielding more discriminative and stable representations.
Figure 2: (a) Dog/cat images, despite lacking visual details, may still be prioritized due to strong textual representations, while visually rich samples like a window may be misclassified as the nearest known category, such as a door, due to weak textual alignment. (b) Certain image regions that are visually associated with the term ”picture” may be erroneously emphasized during feature evaluation.
The test-time generalization of BPRE is evaluated across 15 diverse recognition datasets under natural distribution shifts and cross-dataset generalization. As shown in Fig. 1, BPRE demonstrates competitive performance in cross-dataset generalization with both ResNet-50 and ViT-B/16 architectures.
In summary, our contributions can be listed as:
@@
  • We introduce the Multi-Dimensional Quality-Aware Reward Module, integrating Similarity, Confidence, and Diversity Reward Scores for comprehensive feature evaluation, improving prototype selection.
@@
  • We propose the Prototype-Reward Interactive Evolution Module, establishing a bidirectional reinforcement mechanism where reward scores guide prototype evolution, and refined prototypes enhance reward estimation, improving model robustness.
@@
  • Experiments across 15 diverse image recognition datasets confirm the competitive generalization performance of BPRE.
@@
2 Related Works
2.1 Vision-Language Models
Large-scale VLMs such as CLIP [23] and SigLIP [35] have demonstrated remarkable success in visual understanding tasks by leveraging massive image-text pre-training [6, 39]. Their strong zero-shot transferability makes them effective backbones for tasks like image recognition [10, 19] and object detection [31, 17]. To better adapt VLMs to downstream tasks, researchers have proposed two main categories of methods: prompt learning methods [43, 42, 15, 45, 3] and adapter-based methods [40, 34, 9, 44]. Prompt learning methods like CoOp [43] and CoCoOp [42] focus on learning continuous prompt vectors with few-shot labeled data to bridge the domain gap between pre-training and downstream tasks. These methods achieve impressive performance but require labeled training samples that are often unavailable in real-world scenarios. Despite the success of adapter-based methods like Tip-Adapter [40] and TaskRes [34] in enhancing visual-textual representations, their dependence on labeled target data remains a critical bottleneck. Here we present a test-time adaptation framework that overcomes this limitation by enabling real-time model adaptation without prior exposure to target domain labels or training samples. Our approach facilitates robust model deployment across diverse, unfamiliar environments where labeled data collection is unfeasible.
2.2 Test-Time Adaptation
Test-time adaptation addresses domain shift challenges by enabling dynamic model adaptation during deployment using unlabeled test samples [28, 14, 37]. This approach has proven effective across various computer vision tasks, enhancing model robustness and maintaining performance when encountering out-of-distribution scenarios in real-world applications. Test-time adaptation of VLMs has recently gained significant research attention. TPT [25] pioneered this direction by leveraging consistency across augmented views of test samples. The following works enhanced this approach: DiffTPT [8] introduced diffusion-based augmentations, while C-TPT [33] focused on addressing calibration errors. Moving beyond single-sample adaptation, TDA [14] and DMN [41] introduced memory-based mechanisms to leverage information across multiple test samples, marking a shift from instance-level to collective adaptation strategies. DPE [37] proposes a dual-modality prototype evolution mechanism, which achieves more accurate feature modeling through dynamic updating and optimization of dual-modality feature representations during testing. However, existing methods including DPE [37] focus solely on prototype evolution, overlooking the mutual interaction between prototypes and sample quality. We address this by proposing a Bidirectional Prototype-Reward co-Envolution framework that enables quality-aware prototype evolution while ensuring consistent feature representation through mutual guidance.
3 Method
@@
3.1 Preliminary
Zero-Shot CLIP. CLIP [23] consists of dual encoders: a visual encoder $\mathcal{E}_{v}(\cdot)$ and a text encoder $\mathcal{E}_{t}(\cdot)$, which project images and text into a shared $d$-dimensional embedding space. CLIP performs zero-shot prediction for a $C$-class classification task by measuring similarities between image and text embeddings. Given a test image $X_{\mathtt{test}}\in\mathcal{D}_{\mathtt{test}}$ and class-specific text descriptions ${\mathcal{T}_{c}}$, the prediction process can be formulated as:
$f_{v}=\mathcal{E}_{v}(X_{\mathtt{test}}),\quad f_{t_{c}}=\mathcal{E}_{t}(\mathcal{T}_{c})$
$\mathbb{P}_{CLIP}(y=y_{c}|X_{\mathtt{test}})=\frac{\exp(\mathrm{sim}(f_{t_{c}},f_{v})/\tau)}{\sum_{t^{\prime}}\exp(\mathrm{sim}(f_{t^{\prime}},f_{v})/\tau)},$ (1)
where $\mathrm{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is the temperature parameter.
Test-Time Prompt Tuning. TPT [25] improves CLIP’s zero-shot generalization by learning adaptive prompts from test samples. For each test image $X_{\mathtt{test}}$, TPT generates $N$ augmented views ${\mathcal{A}_{n}(X_{\mathtt{test}})}_{n=1}^{N}$ and aggregates their predictions based on confidence. Specifically, TPT selects predictions with entropy below threshold $t$ and computes their average within the top $\rho$-percentile:
@@ 3.1 Preliminary
×$\quad\times\mathbb{P}_{\mathtt{CLIP}}(\mathcal{A}_{n}(X_{\mathtt{test}})),$ (2)
@@
3.2 Multi-Dimensional Quality-Aware Reward
To effectively assess samples’ contributions to model performance and guide feature learning, we introduce a Multi-Dimensional Quality-Aware Reward Module, which dynamically assigns quality scores by integrating multiple characteristics.
Similarity Reward Score. In cross-modal learning, semantic consistency between sample features and category prototypes serves as a crucial indicator for sample quality evaluation. To this end, we design a similarity-based reward $R_{sim}$ that measures the alignment between sample features and prototype representations:
@@ 3.2 Multi-Dimensional Quality-Aware Reward
Confidence Reward Score. To assess the model’s prediction confidence, we define a confidence reward $R_{conf}$ based on the entropy of the predicted probability distribution. Utilizing temperature-scaled normalization, $R_{conf}$ is derived from CLIP’s predicted probability distribution $\mathbf{P}_{CLIP}$:
@@
Diversity Reward Score. To prevent representation collapse and encourage diverse feature exploration, we introduce a diversity reward $R_{div}$ based on historical feature comparisons. We maintain a fixed-size feature memory queue =$\mathcal{H}=\{h_{1},h_{2},\dots,h_{M}\}$, where each $h_{j}\in\mathbb{R}^{d}$ represents the average feature of a past batch, and $M$ is the queue size. For a given sample feature $f_{i}$, its novelty is measured by the maximum cosine similarity with historical features:
@@
This framework employs a comprehensive reward mechanism to optimize cross-modal transfer learning. By integrating similarity, confidence, and diversity metrics, it dynamically evaluates feature representations. The adaptive reward system, guided by a warmup phase and historical feature tracking, ensures stable learning signals while facilitating prototype evolution, effectively bridging reward estimation and prototype learning.
3.3 Prototype-Reward Interactive Evolution
Building on the Multi-Dimensional Quality-Aware Reward Module, we propose a dynamic prototype evolution mechanism that utilizes reward-guided learning to refine category-specific prototypes. By weighting feature contributions based on their computed rewards, this approach adaptively enhances prototype representations, fostering a more robust and discriminative feature space for cross-modal recognition.
Prototype Representation and Initialization. For a classification task with $C$ categories, we maintain a set of prototype representations in a $d$-dimensional image feature space: $v_{c}\in\mathbb{R}^{d}$. These prototypes act as learnable category-level representations, capturing essential class characteristics.
At the start of training, prototypes are initialized using category-specific statistics. The visual prototype $v_{c}^{(0)}$ for category $c$ is set as the mean feature representation of all samples in that category:
@@ 3.3 Prototype-Reward Interactive Evolution
Quality-Aware Visual Prototype Evolution. Leveraging the comprehensive quality assessment derived from our multi-dimensional reward mechanism, we propose a quality-aware momentum update strategy for dynamic prototype evolution. This mechanism strategically incorporates sample contributions by utilizing their reward scores $R$ as quality indicators, thereby ensuring that samples with higher reward values exert greater influence on the prototype evolution process. Specifically, for category $c$, the visual prototype update at time step $t$ is formulated as:

### v2

@@ Abstract
Test-time adaptation (TTA) is crucial in maintaining performance of Vision Language Models (VLMs) when facing distribution shifts, particularly when the source data or target labels are inaccessible. Existing TTA methods predominantly leverage the output probability distribution of CLIP for feature evaluation, resulting in biases under domain shifts, which cause misclassified features due to text priors or incorrect textual associations. To address these issues, we propose Bidirectional Prototype-Reward co-Evolution (BPRE), a novel VLMs framework with TTA that integrates feature quality assessment with prototype evolution via a synergistic feedback loop. First, the Multi-dimensional Quality-aware Reward Module (MQRM) is designed to evaluate feature quality and guide prototype refinement precisely. The continuous refinement of prototype quality via Prototype-Reward Interactive Evolution (PRIE) enhances the computation more robust. Through this bidirectional interaction, the precision of rewards and prototype evolution mutually reinforce each other, forming a self-evolving feedback cycle. Extensive experiments conducted on 15 diverse recognition datasets demonstrate that our model consistently achieves superior performance compared to other SOTA methods, and advances VLM generalization capabilities through emphasizing comprehensive feature evaluation.
Index Terms:
Prototype Learning, Test-Time Adaptation, Visual Language Models.
I Introduction
Recent advances in multimodal models [1, 2, 3, 4, 5] have opened new opportunities to integrate human language into various computer vision tasks. VLMs such as CLIP [6] demonstrate remarkable zero-shot generalization capabilities, enabling them to perform well on tasks without task-specific training. Their performance drops significantly when faced with domain shifts. This decline is mainly due to discrepancies between the data distributions of the source and target domains. Moreover, such domain discrepancies are inevitable in dynamic and ever-changing environments, making effective domain adaptation for VLMs an urgent and challenging problem.
Figure 1: (a) Images of dogs or cats may lack visual details, but they can still be prioritized due to their strong textual representations. In contrast, visually rich samples, like a window, may be misclassified as the nearest known category, such as a door, if their textual alignment is weak.This happens when their textual alignment is weak. (b) Certain image regions may be visually associated with the term “picture.” These regions can be erroneously emphasized during feature evaluation.
Efficient adaptation methods [7, 8, 9, 10] have been proposed to fine-tune VLMs using only a limited number of target domain samples. The requirement for target domain annotations remains a critical bottleneck, especially in dynamic environments with frequent distribution shifts [11, 12]. In addition, several studies have explored the test-time adaptation (TTA) for VLMs, aiming to efficiently optimize the model during the test phase with minimal cost. Shu et al. [13] proposed test-time prompt tuning (TPT), which optimizes prompts for individual test samples to improve out-of-distribution generalization. This framework was further enhanced by DiffTPT [14], which incorporates diffusion-based augmentations. In addition, Karmanov et al. [15] introduced a training-free approach based on dynamic visual caches. Zhang et al. [16] developed a dual prototype evolving strategy that accumulates task-specific knowledge from both visual and textual modalities during testing.
Most existing TTA methods [7, 15, 16] emphasize leveraging historical feature information to accumulate class-specific task knowledge via caches and prototypes. These approaches construct accurate and discriminative target class representations, which are critical for enhancing model robustness to distribution shifts and improving efficiency during test time. Current methods [15, 16] primarily evaluate feature quality using CLIP’s output probability distribution, which fuses text and image features, and inherently uses the text modality as a reference. However, distribution shifts during testing resulting in misalignment between text and image embedding [17], leading to biases in feature selection and evaluation, as illustrated in Fig. 1. On the one hand, text priors drive the model to prioritize features aligning strongly with pre-trained text descriptions, even when such features lack adequate visual information, while misclassifying visually rich samples with weak text representations (Fig. 1(a)). On the other hand, misalignment of text descriptions introduces feature evaluation biases, where CLIP incorrectly associates generalized text descriptions (e.g.,“picture”) with certain image features, leading the model to mistakenly consider them high-quality samples while overlooking their true semantics, thereby affecting accuracy of feature selection (Fig. 1(b)).
Inspired by these insights, we propose a novel framework for test-time adaptation of VLMs, named Bidirectional Prototype-Reward co-Evolution (BPRE). BPRE iteratively refines task-specific knowledge, as illustrated in Fig. 2(c). In contrast to existing methods, we introduce the Multi-dimensional Quality-aware Reward Module (MQRM), which evaluates features by incorporating intrinsic image properties beyond mere textual correspondence. Integrating visual information with semantic similarity and feature diversity, this module enables comprehensive image feature quality assessment, thereby mitigating excessive reliance on text modality. Semantic similarity ensures that selected features align with the task-specific semantic space, reducing errors caused by text biases. Meanwhile, feature diversity broadens coverage across varied data distributions, preventing overfitting and improving prototype selection. In addition, the Prototype-Reward Interactive Evolution(PRIE) mechanism adaptively updates historical features and Prototypes based on reward scores. This bidirectional reinforcement loop allows refined prototypes to improve reward estimation, creating a self-reinforcing optimization process that boosts noise robustness, promotes interpretable cross-modal learning, and yields more discriminative and stable representations. The test-time generalization of BPRE is evaluated on 15 diverse recognition datasets under natural distribution shifts and cross-dataset scenarios. As shown in Fig. 2(a)(b), BPRE achieves competitive cross-dataset generalization performance with both ResNet-50 and ViT-B/16 architectures.
@@ I Introduction
Our main contributions are summarized as follows:
@@
  • We introduce the Multi-dimensional Quality-aware Reward Module (MQRM), integrating similarity, confidence, and diversity reward scores for comprehensive feature evaluation to improve prototype selection.
@@
  • We propose the Prototype-Reward Interactive Evolution(PRIE) Module, establishing a bidirectional reinforcement process where reward scores guide prototype evolution, and refined prototypes enhance reward estimation.
@@
  • Experiments on 15 diverse image recognition datasets demonstrate that our method achieves competitive generalization performance and advances the generalization capabilities of VLMs.
@@
II Related Works
II-A Vision-Language Models
Large-scale VLMs such as CLIP [6] and SigLIP [18] have demonstrated remarkable success in visual understanding tasks by leveraging massive image-text pre-training [19, 20, 21]. Their strong zero-shot transferability makes them effective backbones for tasks like image recognition [22, 23] and object detection [24, 25]. To better adapt VLMs to downstream tasks, researchers have proposed two main categories of methods: prompt learning methods [26, 27, 28, 8, 9] and adapter-based methods [7, 29, 30, 31]. Prompt learning methods like CoOp [26] and CoCoOp [27] focus on learning continuous prompt vectors with few-shot labeled data to bridge the domain gap between pre-training and downstream tasks. These methods achieve impressive performance but require labeled training samples that are often unavailable in real-world scenarios. Despite the success of adapter-based methods like Tip-Adapter [7] and TaskRes [29] in enhancing visual-textual representations, their dependence on labeled target data remains a critical bottleneck. Here we present a test-time adaptation framework that overcomes this limitation by enabling real-time model adaptation without prior exposure to target domain labels or training samples. Our approach facilitates robust model deployment across diverse, unfamiliar environments where labeled data collection is unfeasible.
II-B Test-Time Adaptation
Test-time adaptation addresses domain shift challenges by enabling dynamic model adaptation during deployment using unlabeled test samples [32, 15, 16, 33]. This approach has proven effective across various computer vision tasks, enhancing model robustness and maintaining performance when encountering out-of-distribution scenarios in real-world applications. Test-time adaptation of VLMs has recently gained significant research attention. TPT [13] pioneered this direction by leveraging consistency across augmented views of test samples. The following works enhanced this approach: DiffTPT [14] introduced diffusion-based augmentations, while C-TPT [34] focused on addressing calibration errors. Moving beyond single-sample adaptation, TDA [15] and DMN [35] introduced memory-based mechanisms to leverage information across multiple test samples, marking a shift from instance-level to collective adaptation strategies. DPE [16] proposes a dual-modality prototype evolution mechanism, which achieves more accurate feature modeling through dynamic updating and optimization of dual-modality feature representations during testing. However, existing methods, including DPE [16], focus solely on prototype evolution, overlooking the mutual interaction between prototypes and sample quality. We address this by proposing a Bidirectional Prototype-Reward co-Envolution framework that enables quality-aware prototype evolution while ensuring consistent feature representation through mutual guidance.
III Method
@@
III-A Preliminary
III-A1 Zero-Shot CLIP
CLIP [6] consists of dual encoders: a visual encoder $\mathcal{E}_{v}(\cdot)$ and a text encoder $\mathcal{E}_{t}(\cdot)$, which project images and text into a shared $d$-dimensional embedding space. CLIP performs zero-shot prediction for a $C$-class classification task by measuring similarities between image and text embeddings. Given a test image $X_{\mathtt{test}}\in\mathcal{D}_{\mathtt{test}}$ and class-specific text descriptions ${\mathcal{T}_{c}}$, the prediction process can be formulated as:
$\mathbb{P}_{\mathrm{CLIP}}(y=y_{c}\mid X_{\mathtt{test}})=\frac{\exp\left(\mathrm{sim}(\mathcal{E}_{t}(\mathcal{T}_{c}),\mathcal{E}_{v}(X_{\mathtt{test}}))/\tau\right)}{\sum_{t^{\prime}}\exp\left(\mathrm{sim}(f_{t^{\prime}},\mathcal{E}_{v}(X_{\mathtt{test}}))/\tau\right)},$ (1)
where $\mathrm{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is a temperature scaling parameter.
III-A2 Test-Time Prompt Tuning
TPT [13] improves CLIP’s zero-shot generalization by learning adaptive prompts from test samples. For each test image $X_{\mathtt{test}}$, TPT generates $N$ augmented views ${\mathcal{A}_{n}(X_{\mathtt{test}})}_{n=1}^{N}$ and aggregates their predictions based on confidence. Specifically, TPT selects predictions with entropy below threshold $t$ and computes their average within the top $\rho$-percentile:
@@ III-A2 Test-Time Prompt Tuning
×$\quad\times\mathbb{P}_{\mathtt{CLIP}}(\mathcal{A}_{n}(X_{\mathtt{test}}))$ (2)
@@
III-B Multi-dimensional Quality-aware Reward Module
To effectively assess samples’ contributions to model performance and guide feature learning, we introduce a Multi-dimensional Quality-aware Reward Module(MQRM), which dynamically assigns quality scores by integrating multiple characteristics.
III-B1 Similarity Reward Score
In cross-modal learning, semantic consistency between sample features and category prototypes serves as a crucial indicator for sample quality evaluation. To this end, we design a similarity-based reward $R_{sim}$ that measures the alignment between sample features and prototype representations:
@@ III-B1 Similarity Reward Score
III-B2 Confidence Reward Score
To assess the model’s prediction confidence, we define a confidence reward $R_{conf}$ based on the entropy of the predicted probability distribution. Utilizing temperature-scaled normalization, $R_{conf}$ is derived from CLIP’s predicted probability distribution $\mathbf{P}_{CLIP}$:
@@ III-B2 Confidence Reward Score
III-B3 Diversity Reward Score
To prevent representation collapse and encourage diverse feature exploration, we introduce a diversity reward $R_{div}$ based on historical feature comparisons. We maintain a fixed-size feature memory queue =$\mathcal{H}=\{h_{1},h_{2},\dots,h_{M}\}$, where each $h_{j}\in\mathbb{R}^{d}$ represents the average feature of a past batch, and $M$ is the queue size. For a given sample feature $f_{i}$, its novelty is measured by the maximum cosine similarity with historical features:
@@ III-B3 Diversity Reward Score
III-C Prototype-Reward Interactive Evolution
Building on our MQRM, we propose a dynamic prototype evolution mechanism that weights feature contributions based on their computed rewards. This approach adaptively refines category-specific prototypes, creating a more robust and discriminative feature space for cross-modal recognition.
III-C1 Prototype Representation and Initialization
For a $C$-category classification task, we maintain prototype representations $v_{c}\in\mathbb{R}^{d}$ in the image feature space. Initially, the visual prototype $v_{c}^{(0)}$ for category $c$ is set as:
@@ III-C1 Prototype Representation and Initialization
TABLE I: Robustness Evaluation under Natural Distribution Shifts. We report top-1 accuracy (%) across methods using CLIP’s ResNet-50 and ViT-B/16 backbones. Bold and underlined values indicate the best and second-best performance, respectively.
Method Source ImageNet ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Average OOD Average
CLIP-ResNet-50 [6] ICML’21 58.16 21.83 51.41 56.15 33.37 44.18 40.69
CoOp [26] IJCV’22 63.33 23.06 55.40 56.60 34.67 46.61 42.43
CoCoOp [27] CVPR’22 62.81 23.32 55.72 57.74 34.48 46.81 42.82
Tip-Adapter [7] ECCV’22 62.03 23.13 53.97 60.35 35.74 47.04 43.30
TPT [13] NeurIPS’22 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT [14] ICCV’23 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA [15] CVPR’24 61.35 30.29 55.54 62.58 38.12 49.58 46.63
DMN-ZS [35] CVPR’24 63.87 28.57 56.12 61.44 39.84 49.97 46.49
DPE [16] NeurIPS’24 63.41 30.15 56.72 63.72 40.03 50.81 47.66
TPS [38] WACV’25 61.47 30.48 54.96 62.87 37.14 49.38 46.36
BPRE Ours 63.77 30.77 56.97 63.94 39.91 51.07 47.90
CLIP-ViT-B/16 [6] ICML’21 66.73 47.87 60.86 73.98 46.09 59.11 57.20
CoOp [26] IJCV’22 71.51 49.71 64.20 75.21 47.99 61.72 59.28
CoCoOp [27] CVPR’22 71.02 50.63 64.07 76.18 48.75 62.13 59.91
Tip-Adapter [7] ECCV’22 70.75 51.04 63.41 77.76 48.88 62.37 60.27
TPT [13] NeurIPS’22 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT [14] ICCV’23 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA [15] CVPR’24 69.51 60.11 64.67 80.24 50.54 65.01 63.89
DMN-ZS [35] CVPR’24 72.25 58.28 65.17 78.55 53.20 65.49 63.80
DPE [16] NeurIPS’24 71.91 59.63 65.44 80.40 52.26 65.93 64.43
TPS [38] WACV’25 70.19 60.08 64.73 80.27 49.95 65.04 63.76
BPRE Ours 72.22 60.67 65.50 80.68 53.45 66.50 65.08
III-C2 Quality-Aware Visual Prototype Evolution
Leveraging the comprehensive quality assessment derived from our multi-dimensional reward mechanism, we propose a quality-aware momentum update strategy for dynamic prototype evolution. This mechanism strategically incorporates sample contributions by utilizing their reward scores $R$ as quality indicators, thereby ensuring that samples with higher reward values exert greater influence on the prototype evolution process. Specifically, for category $c$, the visual prototype update at time step $t$ is formulated as: